OpenStudy (anonymous):

Out of these numbers list the ones that are Natural Numbers: -8, -2/5, 0, 0.4, -8, 0, @RT{5},

OpenStudy (anonymous):

\[-8, \ - \frac { 2 }{ 5 }, 0, 0.4, \sqrt{5}, \sqrt{49}\]

OpenStudy (anonymous):

Look at the equation, not the question I typed out

OpenStudy (anonymous):

@ganeshie8 @geerky42

OpenStudy (anonymous):

Natural numbers are the counting numbers. Pretend you are counting on your fingers. Which of these numbers are counting numbers?

OpenStudy (anonymous):

So would this be my answer: \[-8, 0, 0.4, \sqrt{5}, \sqrt{49}\] in that correct order?

OpenStudy (anonymous):

@ospreytriple

OpenStudy (anonymous):

You can't count -8 on your fingers, or 0.4, etc. Imagine you've got a basket of apples and you're counting them. Those numbers are the natural numbers. {1, 2, 3, ...}

OpenStudy (anonymous):

So my natural numbers would be this answer then perhaps right? \[0, \sqrt{5}, \sqrt{49}\]

OpenStudy (anonymous):

Unfortunately, there is no agreement as to whether zero is a natural number, but my experience has been that it is normally included as well.

OpenStudy (anonymous):

Can you hold up \(\sqrt{5}\) fingers?

OpenStudy (anonymous):

Oh so those are imaginary numbers right?

OpenStudy (anonymous):

No. Imaginary numbers are something else. \(\sqrt{5}\) is what's called an irrational number. So what's your final answer?

OpenStudy (anonymous):

So my only real natural number could be \[0\]. Because it cant be a fracton either correct?

OpenStudy (anonymous):

Well. Do you what \(\sqrt{49}\) equals?

OpenStudy (anonymous):

7?

OpenStudy (anonymous):

Right. So is that a natural number?

OpenStudy (anonymous):

Yes it is. So my natural numbers are: \[0, \sqrt{49}\]?

OpenStudy (anonymous):

Correct. Well done. The set of natural numbers doesn't have any fractions, or decimals, or negative number. Just counting numbers.

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

You're welcome